OpenStudy (olyalhic):

(5x - 4w)(3x +8w) explain please

OpenStudy (deerhunter15):

15x^2-32w^2+28wx

Mehek (mehek14):

Did you learn FOIL @Olyalhic ?

OpenStudy (olyalhic):

@Mehek14 I did a long time ago, so i can't quite remember

OpenStudy (deerhunter15):

did i get it right?

Mehek (mehek14):



OpenStudy (olyalhic):

oh okay so it would be 32w^2 +28wx +15x^2 ?

OpenStudy (olyalhic):

@Mehek14

Mehek (mehek14):

\(\bf{5x*3x\\5x*8w\\-4w*3x\\-4w*8w}\)

Mehek (mehek14):

multiply all of those ^

OpenStudy (olyalhic):

i got 15x,40wx,12wx and 32w. so 15x - 28wx + 32w?

Mehek (mehek14):

\(\bf{5x*3x=15x^2\\5x*8w=40xw\\-4w*3x=-12xw\\-4w*8w=-32w^2}\)

Mehek (mehek14):

add like terms 40xw + (-12xw)

OpenStudy (olyalhic):

oh i see what i missed, \[15x^{2} -32w ^{2}+28xw ^{2}\] is that right?

Mehek (mehek14):

\(\tt{28xw}\) not \(\tt{28xw^2}\)

OpenStudy (olyalhic):

okay i wasn't sure about that thank you <3

Mehek (mehek14):

yw \(\ddot\smile\)